OpenStudy (anonymous):

area shared between the 2 circles r = 6 cos(theta), r = 6 sin(theta)

OpenStudy (anonymous):

okokok

OpenStudy (anonymous):

wat grade r u in

OpenStudy (anonymous):

well

OpenStudy (anonymous):

college calc 3 but doing calc 2 review but I took calc 2 years ago and forgot everything

OpenStudy (anonymous):

oh got u

OpenStudy (anonymous):

wats ur queestion

OpenStudy (anonymous):

sooooooooooooooo

OpenStudy (anonymous):

I need the area shared between the 2 circles, r = 6 cos(theta) and r = 6 sin(theta) , and I am having a hard time determing the interval that is the shared area , I believe you set the 2 equations equal to each other like this, 6cos(theta) = 6sin(theta) ,

OpenStudy (anonymous):

yes that is write than y do u hv this question here

OpenStudy (anonymous):

well................................................

OpenStudy (anonymous):

is it pi/4?

OpenStudy (anonymous):

yes yes yes yes

OpenStudy (anonymous):

im a 7th grader and i know this very well and im a strat a studenet

OpenStudy (anonymous):

maybe you should invest a little time in grammar then

OpenStudy (dumbcow):

@dany10 , do you have the answer?...i am also having trouble setting up polar integral but i solved area using rectangular coord

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

oh yes i know how to do it i had told him

OpenStudy (dumbcow):

@vk278 , haha if you are in 7th grade and can do calc2 you my friend are a genius ...im guessing english is not your first language

OpenStudy (anonymous):

oh no that is not my first lauguage lol

OpenStudy (anonymous):

well

OpenStudy (anonymous):

well I came up with this A =(1/2) intergral from -pi/4 to pi/4 (6cos(theta)^2) - intergral from -pi/4 to pi/4 (6sin(theta)^2 )

OpenStudy (anonymous):

that is apsullutlly right air five high air five i mean

OpenStudy (anonymous):

@vk278 I think your just trolling lol and don't know what i am taking about

OpenStudy (anonymous):

i do ido lol

OpenStudy (anonymous):

so any other questions i could answer for u

OpenStudy (anonymous):

welllllllllllll................................

OpenStudy (dumbcow):

ok your limits are off, just figured it out  1. \[\frac{1}{2} \int\limits _0 ^{\pi/4} (6\sin \theta)^{2} d \theta\] 2. \[\frac{1}{2} \int\limits\limits _{\pi/4} ^{\pi/2} (6\cos \theta)^{2} d \theta\] add them together to get total area